# DAG-Plan: Generating Directed Acyclic Dependency Graphs for Dual-Arm Cooperative Planning

### v2

@@ Abstract
Dual-arm robots offer enhanced versatility and efficiency over single-arm counterparts by enabling concurrent manipulation of multiple objects or cooperative execution of tasks using both arms. However, effectively coordinating the two arms for complex long-horizon tasks remains a significant challenge. Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations, failing to fully leverage the capabilities of dual-arm systems. To address this limitation, we introduce DAG-Plan, a structured task planning framework tailored for dual-arm robots. DAG-Plan harnesses large language models (LLMs) to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG). Critically, DAG-Plan dynamically assigns these sub-tasks to the appropriate arm based on real-time environmental observations, enabling parallel and adaptive execution. We evaluate DAG-Plan on the novel Dual-Arm Kitchen Benchmark, comprising 9 sequential tasks with 78 sub-tasks and 26 objects. Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate plans, achieving nearly 50% higher efficiency compared to the single-arm task planning baseline and nearly double the success rate of the dual-arm task planning baseline.
Keywords: Dual-arm Robots, Task Planning, Large Language Models
1 Introduction
@@
Achieving effective bimanual coordination in robotics is challenging due to the complexities of dual-arm operations, requiring precise spatial and temporal coordination [1, 2]. While humans effortlessly coordinate their hands in daily tasks, replicating such coordination in robots presents significant challenges for both planning and learning-based methods. Traditionally, planning-based methods have focused on motion planning, employing hand-designed primitives to manage the movements of two robotic arms [3, 4]. However, these methods often fall short in dynamic or intricate environments as they lack the flexibility required for adaptive task execution. In contrast, learning-based strategies, such as Reinforcement Learning (RL) and Imitation Learning (IL), provide more adaptive solutions by enabling robots to learn control policies from either human-designed rewards [5, 6] or human demonstrations [7, 8]. However, these methods struggle to generalize to zero-shot scenarios, where robots need to execute tasks without prior specific training.
@@
Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [9, 10, 11, 12]. By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to new scenarios and tasks [13, 14, 15, 16, 17, 18]. However, existing planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills. While some studies [19, 20] have employed dual-arm robots as test platforms, these still engage only one arm at a time or rely on predefined bimanual tasks, leading to inefficiencies. In this work, we propose the development of dual-arm cooperative planning using LLMs. This approach facilitates parallel task execution, thereby significantly improving efficiency. Directly generating dual-arm planning schemes as subgoal sequences with LLMs faces significant challenges, primarily two gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex. Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution.
@@
Our main contributions are summarized as follows: 1) We present DAG-Plan, an efficient cooperative task planning framework for mobile dual-arm robots. This framework represents decomposed sub-tasks as a DAG and dynamically assigns these sub-tasks to the appropriate arm based on the real-time environment state; 2) We conduct the Dual-arm Kitchen Benchmark based on Sapien engine [21], which consists of 9 long-horizon tasks with 78 sub-tasks and 26 assets and objects. This Benchmark is manually constructed by robotic experts and includes plan tests and physical simulation tests; 3) Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods. DAG-Plan achieves nearly a 50% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan. Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates nearly double the success rate.
Figure 1: An overview of DAG-Plan. The DAG-Plan generates a DAG based on human instruction and environmental description. It checks the graph’s completeness and reflects the LLM to regenerate if incomplete. Once a valid DAG is obtained, DAG-Plan performs task inference to identify executable candidate nodes. The occupied arm and free arm are assigned priority candidate nodes and common candidate nodes respectively. The framework then evaluates all candidate combinations for feasibility and cost. DAG-Plan selects the nodes with the lowest cost and employs motion planning and reinforcement learning for execution. DAG-Plan updates the graph, iterating inference until the DAG is fully executed.
2 Related Works
@@
Task Planning with LLMs. LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language. Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [22, 23, 19, 24, 25, 20, 26, 27, 28]. However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep. Although LLM+P [19] and CoPAL [20] both have designed the tasks for dual-arm robots. The plan generated by LLM+P and CoPAL still can not manipulate two different objects with each arm at the same time causing low execution efficiency. Additionally, these studies [19, 27, 28] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability.
@@
Dual-arm Robot Manipulation. Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [29, 30] with fixed process operations and domestic settings [31, 32] with single-skill operations. With the development of methods such as motion planning [30, 32], reinforcement learning [33, 34], and imitation learning [35, 36], dual-arm robots can perform many human-like operations at the skill level. However, they still lack the ability to autonomously plan in zero-shot complex scenarios. The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously plan in zero-shot complex scenarios.
@@
Structured Task Decomposition (STD). STD involves breaking down a complex task into a DAG. Previous methods for STD, such as Crowd-Sourced STD [37, 38] and Query-based STD [39, 40], were limited by data availability. However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD. TaskLAMA [41] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies. In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
3 Method
@@
We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node. The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed. This ensures that dependencies are meticulously adhered to, allowing for efficient task execution. The robot dynamically assesses the state of its environment and the status of its arms to select the next optimal tasks from the graph. This ongoing selection process prioritizes activities, keeping both arms continuously engaged, either cooperatively or independently, depending on task requirements. By emphasizing the directed nature of these task sequences, the methodology enhances operational efficiency, minimizes idle time, and ensures a smooth workflow, significantly improving the robot’s performance in complex operational environments. An overview of the DAG-Plan pipeline is illustrated in Figure 1.
3.1 Directed Acyclic Sub-task Dependency Graph Generation
@@
To address this limitation, we propose a novel approach using LLMs to generate optimized dual-arm plans. Our method involves decomposing complex tasks into a dual-arm task graph instead of a linear sequence. This graph better represents the complex temporal dependencies of bimanual operations. We define the graph as $G=(V,E,T,N)$, where $V$ denotes the tasks, $E$ represents the dependencies, $T$ categorizes the task types, and $N$ specifies the arm requirements. Each vertex $v_{i}\in V$ corresponds to a specific sub-task, and each directed edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that $v_{i}$ must be completed before $v_{j}$. The task types include: 1) Occupy, where tasks involve the engagement of the robot’s gripper and the arm will be occupied after execution, typically for grasping or holding an object; 2) Tool use, which refers to tasks that require the use of a tool, remaining in the gripper throughout the operation; 3) Release, for tasks where an object is released from the gripper, often associated with placement or release into a specific location; 4) Operate, denoting general operational tasks that leave the gripper free post-completion, where the arm will be occupied during execution and released afterward; and 5) Complete, which marks the end of all tasks, represented as the terminal node in the graph. This classification aids in specifying the nature of the task and the number of arms required, thereby enabling a more sophisticated and efficient planning strategy tailored for dual-arm robotic systems. The arm number of node $n_{i}\in N$ represents the arm number of node needed. If $n_{i}=1$ a node requires one arm. If $n_{i}=2$, both arms are needed for execution. The occupy-release pairs are crucial structures in dual-arm task graphs. An occupy-release pair mainly consists of an occupy node as start point and a release node as end point, with potentially several tool use nodes in between. A complete occupy-release pair ensures that the robot arm is not continuously occupied and prevents placing an object without first grasping it. Additionally, the dual-arm task graph should be a fully connected DAG. After generating the dual-arm task graph, we check it for completeness. If the graph contains incomplete grasp-release pairs or is not fully connected, we reflect the LLMs to regenerate the task graph.
Figure 2: The process of Task Planning Inference. In the task “cut the carrot and red bell pepper", DAG-Plan initializes common candidate nodes based on the DAG. It evaluates node combinations, checks feasibility, and calculates costs. The right arm is selected to grasp the carrot and the left to grasp the red bell pepper. After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for each arm. In stage 2, each arms are assigned corresponding priority candidate nodes, checked, and executed. The task graph and nodes are updated again. The priority nodes become empty, indicating the arms are free. In stage 3, with only a single node left, the closest arm to the target executes this node.
@@
3.2 Task Planning Inference with Generated Directed Acyclic Graph
@@
After generating the dual-arm task graph, the planning process enters the inference phase. This phase utilizes the task graph and the observed state to dynamically refine the planning of robotic arm operations. It selects and executes sub-tasks that are executable and have the lowest cost, based on the task graph and the current environment state. As shown in Figure 2, we provide a detailed and specific illustration of the task planning inference process. We first identify the two types of candidate nodes in DAG-Plan: common candidate nodes and priority candidate nodes. The common candidate nodes include those that can be executed when the robotic arm is in the free state. The priority candidate nodes include subsequent nodes when the arm is already engaged in an occupy-release pair. The common candidate nodes are initially selected by identifying nodes within the graph that no other nodes point to.
@@
Once we obtain the candidate nodes for each robotic arm, we generate all possible combinations of left and right arm candidate nodes. These combinations are then checked for feasibility, and any pairs that fail the checks are removed from the candidate set. There are three checks in total. The first check involves verifying the presence of an occupy node within the candidates. If an occupy node is found, we further examine whether its successor nodes contain dependencies that require other conditions to be met first. This could lead to prolonged arm occupancy, thereby decreasing operational efficiency. The second check assesses the distance between target objects for the left and right arms. If these targets are beyond a specified distance threshold, it becomes difficult for both arms to operate simultaneously. The third check evaluates the relative positions of the target locations for each arm. If the target position for the left arm is to the right of the right arm’s target, there is a risk of the arms crossing and colliding, which would prevent the execution of these candidate nodes. After completing the checks, we calculate the cost of the left and right hand candidate nodes based on the environment state. We aim for the target objects to be close to the robotic arms and to each other, facilitating dual-arm operations. Therefore, the cost is composed of two parts: the distance from each target to the respective arm, and the distance between the targets. The pair with the lowest combined cost is selected for execution. The cost $J$ is represented as:
$J=\text{dis}(obj_{\text{right}},obj_{\text{left}})+\frac{1}{2}\left[\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}})\right].$
3.3 Motion Planning and Reinforcement Learning Mixed Skill Learning
Once the sub-tasks to be executed are determined, the robot needs to perform the corresponding actions to bridge the gap between textual instructions and the physical environment. Reinforcement Learning (RL) is a powerful technique that enables robots to acquire and refine skills autonomously by interacting with their environment[42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. By engaging in a trial-and-error process and receiving feedback in the form of rewards, robots can learn to optimize their actions to achieve specific objectives [55, 56, 57, 58, 59, 60]. Recent advancements [61, 62, 63, 64] have further enhanced automation by using LLMs to generate text-to-reward mappings. We combine motion planning (RRT-connect) [65] and RL (PPO) [47] to facilitate skill learning (see details in Appendix A). The robot approaches the target object using motion planning and then learns the corresponding skill through RL. We input the generated dual-arm task graph and description of the reward APIs into LLMs, which then combine different reward APIs and determine if they are necessary conditions for completing the task, thus achieving skill learning. However, given the sensitivity of RL training and the ambiguity of language, RL policies may fail to achieve the goal or achieve it in unexpected ways. Therefore, we manually fine-tune the generated rewards to ensure that the final skills meet the task requirements.
4 Experiments
4.1 Experimental Setup

### v3

@@ Abstract
Dual-arm robots offer enhanced versatility and efficiency over single-arm counterparts by enabling concurrent manipulation of multiple objects or cooperative execution of tasks using both arms. However, the coordination of dual-arm systems for long-horizon tasks continues to pose significant challenges, stemming from the intricate temporal and spatial dependencies among sub-tasks, necessitating intelligent decisions regarding the allocation of actions between arms and their optimal execution order. Existing task planning methods predominantly focus on single-arm robots or rely on predefined bimanual operations to use large language models (LLMs) generate task sequence with linear temporal dependency, failing to fully leverage the capabilities of dual-arm systems. To address this limitation, we introduce DAG-Plan, a structured task planning framework tailored for dual-arm robots. DAG-Plan harnesses LLMs to decompose intricate tasks into actionable sub-tasks represented as nodes within a directed acyclic graph (DAG). Critically, DAG-Plan dynamically assigns these sub-tasks to the appropriate arm based on real-time environmental observations, enabling parallel and adaptive execution. We evaluate DAG-Plan on the Dual-Arm Kitchen Benchmark, comprising 5 sequential tasks with 44 sub-tasks. Extensive experiments demonstrate the superiority of DAG-Plan over directly using LLM to generate linear task sequence, achieving 52.8% higher efficiency compared to the single-arm task planning and 48% higher success rate of the dual-arm task planning. Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time. More demos and information are available on https://sites.google.com/view/dag-plan.
Index Terms:
Dual-arm Robots, Task Planning, LLMs
I Introduction
@@
Achieving effective bimanual coordination in robotics is challenging due to the complexities and long-horizon of dual-arm operations, requiring precise spatial and temporal coordination [1, 2]. While humans effortlessly coordinate their hands in daily long-horizon tasks, replicating such coordination in robots presents significant challenges. Traditional method employed hand-designed primitives to manage the movements of dual robotic arms [3, 4]. Whereas these methods often fall short in long-horizon task as they lack the flexibility required for adaptive task execution. Large language models (LLMs) have emerged as powerful tools endowed with extensive knowledge and sophisticated reasoning abilities [5, 6]. By systematically breaking down tasks into actionable sub-tasks and leveraging their commonsense knowledge and implicit reasoning capabilities, LLMs empower robots to effectively adapt to long-horizon scenarios and tasks in the wild [7, 8].
Existing LLMs planning methods are primarily applied to single-arm robots, focusing on using one arm to perform skills. While some studies [9, 10] have employed dual-arm robots as test platforms, these still engage only one arm at a time or rely on predefined bimanual tasks, leading to inefficiencies. Several multi-agent planning methods  [11, 12, 13] have been proposed, which assign tasks to each robot through independent task decomposition or interactions among multiple robots, significantly enhancing performance in multi-robot collaborative tasks. These methods share a common characteristic: they all utilize LLMs to generate task sequences with linear temporal dependencies.
@@
Generating dual-arm planning schemes as task sequences with LLMs faces significant challenges, primarily three gaps: 1) Dual-arm collaboration allows multiple sub-tasks to be executed simultaneously, making the temporal dependencies between them highly complex. Previous approaches, which employed linear temporal dependency lists for task sequencing, have proven to be inefficient for planning and execution; 2) The non-interactivity with the environment, as the execution order of the task sequence and the side of the executing arm are fixed in non-iterative method, making it impossible to choose executable and cost-effective sub-tasks based on the environment and robot state during execution; 3) The iterative approach, while capable of interacting with the environment and achieving a higher success rate compared to non-iterative methods, necessitates querying the LLM for each decision, resulting in significant token consumption and time costs.
@@
Our main contributions are summarized as follows:
  • We identify the limitations of planning methods which generate linear task sequences when applied to dual-arm robots, particularly their inefficiency in handling complex temporal dependencies and their inability to adaptively interact with the environment during task execution.
  • We present DAG-Plan, an efficient cooperative task planning framework for dual-arm robots. This framework represents decomposed sub-tasks as a DAG with LLMs and dynamically assigns sub-tasks to the appropriate arm based on the real-time environment state.
  • Extensive experiments on the Dual-arm Kitchen Benchmark show that DAG-Plan significantly outperforms other methods. DAG-Plan achieves 52.8% increase in efficiency over the baseline by directly employing LLMs to generate the single-arm plan. Compared to the baseline, which directly utilizes LLMs for dual-arm plans, DAG-Plan demonstrates 48% higher success rate. Compared to iterative methods, DAG-Plan improving execution efficiency 84.1% due to its fewer query time.
II Related Works
II-A Task Planning with LLMs
@@
LLMs are increasingly being used to generate sequences of executable actions that enable an agent to achieve goals represented in natural language. Previous studies have successfully utilized the commonsense and in-context learning capabilities of pre-trained LLMs to create executable plans for embodied agents [14, 9, 10, 15, 11, 12, 13]. However, most of these studies have been applied to single-arm robots, only requiring consideration of executing a single-arm action at one timestep. Additionally, these studies [9, 11] explicitly use the operational rules of the environment as input, whereas we only provide an environment description, relying on the LLM’s inherent world modeling capability. Multi-agents planning is the closest to dual-arm planning. Twostep [11] decomposes task into two independent tasks. RoCo [12] and DABICO [13] assign each robot with a LLM to discuss and collectively reason task strategies. These methods either fail to consider the constraints of dual-arm robot or are limited by the linear dependencies of task sequence.
II-B Dual-arm Robot Manipulation
@@
Dual-arm coordination has made progress in industrial [1, 2] and agricultural scenarios [16, 17] with fixed process operations and domestic settings [18, 19] with single-skill operations. With the development of methods such as motion planning [17, 19], foundation model [20, 21], reinforcement learning [22, 23], and imitation learning [24, 25], dual-arm robots can perform many human-like operations at the skill level. However, they still lack the ability to autonomously execution in long-horizon task. The commonsense and contextual learning capabilities of pre-trained LLMs make it possible for dual-arm robots to autonomously execution.
II-C Structured Task Decomposition (STD)
@@
STD involves breaking down a complex task into a DAG. Previous methods for STD, such as Crowd-Sourced STD [26, 27] and Query-based STD [28, 29], were limited by data availability. However, LLMs contain extensive real-world commonsense knowledge, offering new approaches for STD. TaskLAMA [30] has conducted detailed research on structured task decomposition using LLMs, demonstrating that LLMs can decompose real-world tasks into task graphs with temporal dependencies. In this work, we explore the use of DAG to address issues in dual-arm robot sequential planning with low execution efficiency.
III Methods
@@
We utilize LLMs to generate a DAG, where each task for a dual-arm robot is represented as a node. The directed edges between these nodes are crucial as they establish a clear and mandatory sequence of tasks, dictating the order in which tasks must be performed. This ensures that dependencies are meticulously adhered to, allowing for efficient task execution. The robot dynamically assesses the state of its environment and the status of its arms to select next optimal tasks from the graph. This ongoing selection process prioritizes activities, keeping both arms continuously engaged. An overview of DAG-Plan pipeline is illustrated in Figure 1.
Figure 1: An overview of DAG-Plan. The DAG-Plan generates a DAG based on human instruction and environmental description. It checks the graph’s completeness and reflects the LLM to regenerate if incomplete. Once a valid DAG is obtained, DAG-Plan performs task inference to identify executable candidate nodes. The occupied arm and free arm are assigned priority candidate nodes and common candidate nodes respectively. The framework then evaluates all candidate combinations for feasibility and cost. DAG-Plan selects the nodes with the lowest cost and employs skill in library for execution. DAG-Plan updates the graph, iterating inference until the DAG is fully executed.
III-A Directed Acyclic Sub-task Dependency Graph Generation
@@
To address this limitation, we propose a novel approach using LLMs to generate optimized dual-arm plans. Our method involves decomposing complex tasks into a dual-arm task graph instead of a linear sequence. This graph better represents the complex temporal dependencies of bimanual operations. We define the graph as $G=(V,E,T,N)$, where $V$ denotes the tasks, $E$ represents the dependencies, $T$ categorizes the task types, and $N$ specifies the arm requirements. Each vertex $v_{i}\in V$ corresponds to a specific sub-task, and each directed edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that $v_{i}$ must be completed before $v_{j}$. The task types include: 1) Occupy, where tasks involve the engagement of the robot’s gripper and the arm will be occupied after execution, typically for grasping or holding an object; 2) Tool use, which refers to tasks that require the use of a tool, remaining in the gripper throughout the operation; 3) Release, for tasks where an object is released from the gripper, often associated with placement or release into a specific location; 4) Operate, denoting general operational tasks that leave the gripper free post-completion, where the arm will be occupied during execution and released afterward; and 5) Complete, which marks the end of all tasks, represented as the terminal node in the graph. This classification aids in specifying the nature of the task and the number of arms required, thereby enabling a more sophisticated and efficient planning strategy tailored for dual-arm robotic systems. The arm number of node $n_{i}\in N$ represents the arm number of node needed. The occupy-release pairs are crucial structures in dual-arm task graphs. An occupy-release pair mainly consists of an occupy node as start point and a release node as end point, with potentially several tool use nodes in between. A complete occupy-release pair ensures that the robot arm is not continuously occupied and prevents placing an object without first grasping it. Additionally, the dual-arm task graph should be a fully connected DAG. After generating the dual-arm task graph, we check it for completeness. If the graph contains incomplete grasp-release pairs or is not fully connected, we reflect the LLMs to regenerate the task graph.
@@
III-B Task Planning Inference with Generated Directed Acyclic Graph
@@
After generating the dual-arm task graph, the planning process enters the inference phase. This phase utilizes the task graph and the observed state to dynamically refine the planning of robotic arm operations. It selects and executes sub-tasks that are executable and have the lowest cost, based on the task graph and the current environment state. As shown in Figure 2, we provide a detailed and specific illustration of the task planning inference process. We first identify the two types of candidate nodes in DAG-Plan: common candidate nodes and priority candidate nodes. The common candidate nodes include those that can be executed when the robotic arm is in the free state. The priority candidate nodes include subsequent nodes when the arm is already engaged in an occupy-release pair. The common candidate nodes are initially selected by identifying nodes within graph that no other nodes point to.
@@
Figure 2: The process of Task Planning Inference. In task 2 “clean the table (Hard)”, DAG-Plan initializes common candidate nodes based on the DAG. It evaluates node combinations, checks feasibility, and calculates costs. The right arm is selected to grasp apple and the left to open drawer. After execution, the task graph and nodes are updated, adding subsequent release nodes to the priority candidate nodes for right arm. In stage 2, right arm is assigned corresponding priority candidate nodes, checked and left arm still selected node in common candidate nodes with empty priority candidate nodes. The task graph and nodes are updated again. In stage 3, left arm put mug into drawer and right arm grasp lemon.
@@
Once we obtain the candidate nodes for each robotic arm, we generate all possible combinations of left and right arm candidate nodes. These combinations are then checked for feasibility, and any pairs that fail the checks are removed from the candidate set. There are three checks in total. The first check involves verifying the presence of an occupy node within the candidates. If an occupy node is found, we further examine whether its successor nodes contain dependencies that require other conditions to be met first. This could lead to prolonged arm occupancy even dead-lock, thereby decreasing operational efficiency. The second check evaluates the distance between the target objects assigned to the left and right arms. If the distance exceeds a predefined threshold, simultaneous operation of both arms becomes infeasible. The third check analyzes the relative positions of the target locations for each arm. If the target position for the left arm is located to the right of the right arm’s target or is too far away, the arms may collide or fail to reach their targets. To acquire the object positions necessary for decision-making, DAG-Plan utilizes Grounding DINO[31] and SAM2[32] to obtain the point clouds corresponding to the specified objects, thereby determining their locations. For specific categories of objects such as bottles and bowls, the category-specific 6D pose estimation model SAR-Net[33] is employed to estimate their poses.
After completing the checks, we calculate the cost of the left and right hand candidate nodes based on the environment state. We aim for the target objects to be close to the robotic hands facilitating dual-arm operations. The pair with the lowest combined cost is selected for execution. The cost $J$ is represented as:
$\footnotesize J=\text{dis}(obj_{\text{right}},hand_{\text{right}})+\text{dis}(obj_{\text{left}},hand_{\text{left}}).$
III-C Sub-tasks Execution with Foundation Model
Once the sub-tasks to be executed are determined, the robot needs to perform the corresponding actions to bridge the gap between textual instructions and the physical environment. The collective perceptual inferences specific to objects, along with physical insights and parameters for manipulation, are methodically organized and converted into a structured format of executable action code.
III-C1 Occupy
We use the pre-trained AnyGrasp[20] to produce a variety of grasp pose proposals.
III-C2 Tool use
We have specifically designed tool-usage skills based on the categories of tools employed, and have provided adjustable parameters to enable the robot to perform flexible tool-manipulating actions.
III-C3 Release
Based on the current grasping position of the object and the 3D bounding box of the target object, we propose an appropriate release pose to ensure that the target object can be accurately placed at the desired location.
III-C4 Operate
We employ the GAPartNet[21] to forecast the physical characteristics of articulated objects. This model works by dividing the point cloud of an articulated object into its constituent rigid parts and subsequently calculating the articulation parameters.
Figure 3: Snapshots of 5 Tasks of Dual-arm Kitchen Benchmark.
TABLE I: Task list of Dual-arm Kitchen Benchmark.
Index Task Name Human Instruction
Task 1 Clean the table (Easy) Clean the table. Put objects into plate.
Task 2 Clean the table (Hard) Clean the table. The fruits should into the plate. The table should be wiped by sponge. The mug and sponge should in drawer.
Task 3 Stack bowls Stack the bowls onto the wooden tray with the green bowl, blue bowl and yellow bowl order.
Task 4 Make cup of coffee Make a cup of coffee. You should add the coffee, water and milk in order. Finally, stir it with the spoon.
Task 5 Boil vegetables Boil vegetables. Pour water and put vegetables into the pot.
IV Experiments
In the experiments, we are interested in answering the following questions about DAG-Plan:
  • $\bullet$
    To what extent can the task-graph approach improve performance compared to the task-sequence method?
  • $\bullet$
    How does the cost of DAG-Plan compare to that of iterative methods?
  • $\bullet$
    How does the performance of DAG-Plan compare to multi-agents methods?
  • $\bullet$
    Can DAG-Plan be deployed in real-world setting?
IV-A Experimental Setup